# Text-to-Decision Agent: Learning Generalist Policies from Natural Language Supervision

### v2

@@ Abstract
RL systems usually tackle generalization by inferring task beliefs from high-quality samples or warmup explorations. The restricted form limits their generality and usability since these supervision signals are expensive and even infeasible to acquire in advance for unseen tasks. Learning directly from the raw text about decision tasks is a promising alternative to leverage a much broader source of supervision. In the paper, we propose Text-to-Decision Agent (T2DA), a simple and scalable framework that supervises generalist policy learning with natural language. We first introduce a generalized world model to encode multi-task decision data into a dynamics-aware embedding space. Then, inspired by CLIP, we predict which textual description goes with which decision embedding, effectively bridging their semantic gap via contrastive language-decision pre-training and aligning the text embeddings to comprehend the environment dynamics. After training the text-conditioned generalist policy, the agent can directly realize zero-shot text-to-decision generation in response to language instructions. Comprehensive experiments on MuJoCo and Meta-World benchmarks show that T2DA facilitates high-capacity zero-shot generalization and outperforms various types of baselines.
@@
1 Introduction
Reinforcement learning (RL) has emerged as an effective mechanism for training autonomous agents to perform complex tasks in interactive environments [1, 2], unleashing its potential across frontier problems including preference optimization [3], diffusion model training [4], and reasoning like in OpenAI o1 and DeepSeek-R1 [5]. Despite its achievements, one of the grand challenges is generalization: building a general-purpose agent capable of handling multiple tasks in response to diverse user commands [6]. Current RL algorithms are usually specialized to a narrowly defined task distribution and struggle with poor generalization to unseen tasks, due to distribution shift and the lack of self-supervised pre-training techniques [7]. Recent studies usually rely on high-quality samples [8] or warmup explorations [9] to infer task-specific information for generalization at test time, while these supervision signals are expensive and even infeasible to acquire in advance for unseen tasks [10]. This inspires us to explore a much broader source of supervision.
Recently, large language models (LLMs), pre-trained on extensive text corpora, can encode a wealth of semantic knowledge with remarkable representation power and transferability [11]. The typical “text-to-text" interface enables task-agnostic architectures to achieve broad generalization with minimal or no domain-specific data [12]. The availability of large-scale text collections for aggregate supervision during model pre-training has revolutionized the language and multi-modal communities in recent years [13]. Naturally, the above advancements highlight a promising question: Could scalable pre-training methods that learn perception from the supervision embedded in natural language lead to a similar leap forward in developing generalist agents in RL?
@@ 1 Introduction
To tackle these challenges, we propose Text-to-Decision Agent (T2DA), a simple and scalable pre-training framework for learning generalist policies via aligning language knowledge with environment dynamics of decision tasks. First, we pre-train a generalized world model to encode multi-task data into dynamics-aware decision embeddings, effectively capturing task-specific environment dynamics. Second, inspired by CLIP [12], we predict which textual description goes with which decision embedding, efficiently bridging their semantic gap via contrastive language-decision pre-training. It distills the world model structure to the text modality and aligns text embeddings to comprehend the environment dynamics. Finally, we deploy T2DA on two mainstream conditional generation architectures, training generalist policies conditioned on aligned text embeddings. During evaluation, natural language is used to reference learned decision perceptions or describe new ones, and the agent can directly realize zero-shot text-to-decision generation according to textual instructions at hand. Extensive experiments show that T2DA transfers nontrivially to downstream tasks, achieves high-capacity zero-shot generalization, and outperforms various types of baselines.
@@
  • We present a generalized world model designed to capture the environment dynamics, facilitating knowledge grounding in the physical world.
@@
  • We propose a simple and scalable framework that learns generalist policies from natural language supervision, and develop scalable implementations with the potential to train decision models at scale: Text-to-Decision Diffuser and Text-to-Decision Transformer.
@@ 2 Related Work
Generalization in RL. Learning generalist policies is a long-standing challenge in RL. Early methods tackle this from the meta-learning perspective, which is differentiated into the optimization-based meta-RL such as MAML [21] and MACAW [22], and context-based meta-RL like PEARL [23], VariBAD [24], CORRO [25], CSRO [9], and UNICORN [26]. Recent studies of in-context RL attempt to leverage the in-context learning capability of transformer architectures to improve RL’s generalization via casting RL as an across-episodic sequential prediction problem with few-shot prompting at test time [27, 28], such as AD [29] and DPT [30]. In general, the above approaches rely on high-quality samples or domain knowledge to infer task beliefs during evaluation, which limits their generality with this restricted form of supervision. This inspires us to explore a much broader source of supervision from natural language.
LLMs for RL. Existing studies on leveraging knowledge of LLMs for RL domains can be categorized into three main threads: LLMs as policies, LLMs as rewards, and language-conditioned policy learning. The first kind directly utilizes LLMs, or vision-language models (VLMs), as the policy model [15, 31] or high-level semantic planner [14]. Most of them focus on tasks with text-only interactions, i.e., performing textual commands given textual observations [32, 33]. Otherwise, additional modules are needed to map the textual command into actions of other modalities [14, 16]. The second kind uses LLMs or VLMs to automate the generation and shaping of reward functions [34, 17, 35, 18] or state representations [36], providing denser information to facilitate RL training. The above studies usually focus on the single-task setting, while we aim to harness LLM knowledge towards generalist policy learning.
@@
Language-conditioned policy. We focus on the third kind that learns a language-conditioned contextual policy, extracting the world knowledge encoded in LLMs to help train generalist decision agents. Previous studies attempt to translate the language to latent embeddings using pre-trained LLMs like BERT and GPT [37], while these embeddings are learned independently from decision tasks and may fail to capture task-related information. Early methods develop rule-based [38] or domain-specific [39] intermediate representations to capture task-related semantics, which could require a laborious design and lack scalability. BC-Z [19] trains a vision-based robotic manipulation system to achieve zero-shot generalization via conditioning an imitation learning policy on pre-trained embeddings of natural language. To improve language grounding, Pang et al. [40] translates natural language to task language via designing a referential game, and Pang et al. [20] uses supervised fine-tuning to enable bidirectional translation between language-described skills and rollout data. Overall, the challenges mainly include the inability to capture environment dynamics, knowledge misalignment due to semantic gaps, and limited model scalability. In the paper, we tackle these challenges to fully harness LLMs for building generalist decision models.
@@ 3.1 Problem Statement
We consider a language-conditioned multi-task offline RL setting. Tasks follow a distribution $M_{k}\!\!=\!\!\langle\mathcal{S},\mathcal{A},\mathcal{T}_{k},\mathcal{R}_{k},\gamma\rangle\!\!\sim\!\!P(M)$, sharing the same state-action spaces while varying in the reward and state transition functions, i.e., environment dynamics. For each training task $k$, our system receives a user-provided natural language supervision $l_{k}$ that describes the task (e.g., open the door), paired with an offline dataset $\mathcal{D}_{k}=\sum_{i}(s_{i}^{k},a_{i}^{k},r_{i}^{k},s_{i}^{{}^{\prime}k})$ collected by arbitrary behavior policies. The agent can only access the offline datasets $\sum_{k}(l_{k},\mathcal{D}_{k})$ to train a generalist policy $\pi(a|s,l)$ to follow language instructions. At test time, natural language is used to reference learned decision perceptions or describe new ones. The agent can perform text-to-decision generation in test environments in a zero-shot manner, given any language instruction $l_{\text{new}}$ at hand. The primary objective is to obtain a highly generalizable policy that achieves strong zero-shot performance on unseen tasks.
@@ 3.2 Dynamics-Aware Decision Embedding
Natural language is a complex and unbounded representation of human instruction, and training policies to harness language supervision is nontrivial. A standard approach is to convert language instructions to simpler latent embeddings using popular pre-trained language models that can draw on a wealth of knowledge learned from copious amounts of text. However, directly applying these models to decision domains could fail to solve user-specific problems, as they are typically learned independently of decision-making tasks. A major challenge is that LLMs trained on text corpora typically lack grounding in the physical world and fail to capture any environment dynamics.
@@
For effective grounding in decision domains, the agent ought to comprehend the environment dynamics of underlying tasks. RL typically learns from active interactions with the outer environment. The world model, i.e., the reward and state transition functions $p(s^{\prime},r|s,a)$, fully characterizes the environment and presents a promising alternative to capturing the environment dynamics of decision tasks. Therefore, we introduce a generalized world model to encode the decision data from different tasks into a dynamics-aware embedding space. We separate the reasoning about the world model into two parts: i) encoding the dynamics-specific information into a latent embedding, and ii) decoding the environment dynamics conditioned on that embedding.
@@
First, we use a trajectory encoder $\phi$ to abstract multi-task decision trajectories into a compact embedding $z$ that captures task-specific dynamics. Specifically, we “tokenize" different elements (i.e., state, action, and reward) in the raw trajectory sequence $\tau=(s_{0},a_{0},r_{0},...,s_{L},a_{L},r_{L})\sim\mathcal{D}_{k}$, by lifting them to a common representation space as
$e_{t}^{s}\!=\!f_{\phi}^{s}(s_{t}),~{}e_{t}^{a}\!=\!f_{\phi}^{a}(a_{t}),~{}e_{t}^{r}\!=\!f_{\phi}^{r}(r_{t}),~{}~{}~{}\forall t\in[0,L],$ (1)
where $f_{\phi}^{s},f_{\phi}^{a},f_{\phi}^{r}$ are element-specific tokenizers. We arrange the tokens in a 1-D sequence of length $3\times L$ as
$\tau_{e}=(e_{0}^{s},~{}e_{0}^{a},~{}e_{0}^{r},~{}...,~{}e_{L}^{s},~{}e_{L}^{a},~{}e_{L}^{r}).$ (2)
@@
Then, we utilize a bi-directional transformer $E_{\phi}$ to extract the dynamics-aware embedding from the input sequence of tokens as $z_{k}=E_{\phi}(\tau_{e})$. The bi-directional structure allows for implicitly capturing the forward and inverse dynamics.
@@
Second, we introduce a decoder $\varphi$ containing a reward model $\mathcal{R}_{\varphi}$ and state transition model $\mathcal{T}_{\varphi}$. The latent embedding is augmented into the input to predict the instant reward $\hat{r}_{t}\!=\!\mathcal{R}_{\varphi}(s_{t},a_{t};z_{k})$ and next state $\hat{s}_{t+1}\!=\!\mathcal{T}_{\varphi}(s_{t},a_{t};z_{k})$. The encoder-decoder pipeline is jointly trained by minimizing the reward and state transition prediction error conditioned on the dynamics-aware decision embedding as
$\mathcal{L(\phi,\varphi)}=\!\!\!\!\!\!\!\!\!\!$ $\mathbb{E}_{\tau\sim\mathcal{D}_{k}}\left[\mathbb{E}_{z_{k}\sim\phi(\tau)}\left[\mathbb{E}_{t}\left[\left(r_{t}-\mathcal{R}_{\varphi}(s_{t},a_{t};z_{k})\right)^{2}\right.\right.\right.$ (3)
$\left.\left.\left.+\left(s_{t+1}-\mathcal{T}_{\varphi}(s_{t},a_{t};z_{k})\right)^{2}\right]\right]\right],~{}~{}~{}\forall k.$
@@
In practice, we randomly sample a trajectory from the dataset as $\tau\!\sim\!\mathcal{D}_{k}$ and use its decision embedding to decode the dynamics of other trajectories in the same dataset as $\tau^{*}\in\mathcal{D}_{k}\backslash\tau$. The reason is that, since the embedding has access to the entire trajectory’s information, using it to decode the same trajectory can lead to deceptive traps during training. After proper training, the trajectory encoder is frozen for the subsequent learning phases.
@@ 3.3 Contrastive Language-Decision Pre-training
To successfully instruct generalist training with human language, we must bridge the semantic gap between text and the “decision modality". Inspired by CLIP [12], we predict which textual description goes with which training task, bridging their semantic gap via contrastive learning. The core idea is to connect the language supervision with corresponding decision embeddings, distilling the world model structure to the text modality and aligning text embeddings to comprehend the environment dynamics.
Formally, given a batch of $N$ (trajectory $\tau$, textual task description $l$) pairs, we attempt to predict which of the $N\times N$ possible (trajectory, text) pairings across a batch actually occurred. We use the pre-trained trajectory encoder in Sec. 3.2 to extract the dynamics-aware decision embedding from the trajectory as $z=\phi(\tau)$. The text encoder is a transformer-based model $\psi$ that could be initialized from one of a wide variety of popular language models such as CLIP [12] or T5 [41]. Then, the text embedding $z_{T}$ is easily derived by tokenizing the textual instruction $l$ and feeding it to the encoder as $z_{T}=\psi(l)$.
@@
$\text{sim}(\tau,l)=\exp(\alpha)\cdot\frac{\phi(\tau)W_{D}\cdot\psi(l)W_{T}}{||\phi(\tau)W_{D}||\cdot||\psi(l)W_{T}||},$ (4)
$\text{sim}(l,\tau)=\exp(\alpha)\cdot\frac{\psi(l)W_{T}\cdot\phi(\tau)W_{D}}{||\psi(l)W_{T}||\cdot||\phi(\tau)W_{D}||},$
where $\alpha$ is a learnable temperature parameter that controls the range of similarities, and $W_{D}$ and $W_{T}$ are learnable projections that map feature representations of each modality to a joint multi-modal embedding space. Then, the decision-to-text $p(\tau)$ and text-to-decision $p(l)$ similarity scores in the batch are calculated as
@@
Let $q(\tau)$ and $q(l)$ denote the ground-truth similarity scores where negative pairs have a probability of 0 and positive pairs have 1. We optimize a symmetric cross-entropy loss over these similarity scores as
$\mathcal{L}(\psi)=\frac{1}{2}\mathbb{E}_{\tau,l}\left[\text{CE}\left(p(\tau),q(\tau)\right)+\text{CE}\left(p(l),q(l)\right)\right],$ (6)
where CE is the cross-entropy between two distributions. To effectively adapt the text encoder to the semantics of decision modality at a lightweight computing cost, we employ low-rank adaptation (LoRA) [42] to fine-tune the text encoder with the above contrastive loss. Experiments in Appendix F present the analysis of its efficiency and superiority compared to full-parameter fine-tuning.

### v3

# Text-to-Decision Agent: Offline Meta-Reinforcement Learning from Natural Language Supervision

@@ Abstract
Offline meta-RL usually tackles generalization by inferring task beliefs from high-quality samples or warmup explorations. The restricted form limits their generality and usability since these supervision signals are expensive and even infeasible to acquire in advance for unseen tasks. Learning directly from the raw text about decision tasks is a promising alternative to leverage a much broader source of supervision. In the paper, we propose Text-to-Decision Agent (T2DA), a simple and scalable framework that supervises offline meta-RL with natural language. We first introduce a generalized world model to encode multi-task decision data into a dynamics-aware embedding space. Then, inspired by CLIP, we predict which textual description goes with which decision embedding, effectively bridging their semantic gap via contrastive language-decision pre-training and aligning the text embeddings to comprehend the environment dynamics. After training the text-conditioned generalist policy, the agent can directly realize zero-shot text-to-decision generation in response to language instructions. Comprehensive experiments on MuJoCo and Meta-World benchmarks show that T2DA facilitates high-capacity zero-shot generalization and outperforms various types of baselines.
1 Introduction
Reinforcement learning (RL) has emerged as an effective mechanism for training autonomous agents to perform complex tasks in interactive environments [1, 2], unleashing its potential across frontier problems including preference optimization [3], diffusion model training [4], and reasoning such as in OpenAI o1 [5] and DeepSeek-R1 [6]. Despite its achievements, one of the grand challenges is generalization: building a general-purpose agent capable of handling multiple tasks in response to diverse user commands [7]. Due to distribution shift and the lack of self-supervised pre-training techniques, RL agents typically struggle with poor generalization to unseen tasks [8]. Offline meta-RL tackles generalization via training on a distribution of offline tasks [9]. However, they usually rely on high-quality samples [10] or warmup explorations [11] to infer task beliefs for generalization at test time, while these supervision signals are expensive and even infeasible to acquire in advance for unseen tasks [12]. This inspires us to explore a much broader source of supervision.
Recently, large language models (LLMs), pre-trained on extensive text corpora, can encode a wealth of semantic knowledge with remarkable representation power and transferability [13]. The typical “text-to-text" interface enables task-agnostic architectures to achieve broad generalization with minimal or no domain-specific data [14]. The availability of large-scale text collections for aggregate supervision during model pre-training has revolutionized the language and multi-modal communities in recent years [15]. Naturally, the above advancements highlight a promising question: Could scalable pre-training methods that learn perception from the supervision embedded in natural language lead to a similar leap forward in developing generalist agents for offline meta-RL?
@@ 1 Introduction
To tackle these challenges, we propose Text-to-Decision Agent (T2DA), a simple and scalable pre-training framework for offline meta-RL via aligning language knowledge with environment dynamics of decision tasks. First, we pre-train a generalized world model to encode multi-task data into dynamics-aware decision embeddings, effectively capturing task-specific environment dynamics. Second, inspired by CLIP [14], we predict which textual description goes with which decision embedding, efficiently bridging their semantic gap via contrastive language-decision pre-training. It distills the world model structure to the text modality and aligns text embeddings to comprehend the environment dynamics. Finally, we deploy T2DA on two mainstream conditional generation architectures, training generalist policies conditioned on aligned text embeddings. During evaluation, natural language is used to reference learned decision perceptions or describe new ones, and the agent can directly realize zero-shot text-to-decision generation according to textual instructions at hand. Extensive experiments show that T2DA transfers nontrivially to downstream tasks, achieves high-capacity zero-shot generalization, and outperforms various types of baselines.
@@
  • We present a generalized world model designed to capture the environment dynamics, facilitating effective grounding of linguistic knowledge in decision domains.
@@
  • We propose a simple and scalable framework that supervises offline meta-RL with natural language, and develop scalable implementations with the potential to train decision models at scale: Text-to-Decision Diffuser and Text-to-Decision Transformer.
@@ 2 Related Work
Offline Meta-RL learns adaptable policies from pre-collected datasets without requiring online environment interaction [9, 23]. It addresses two fundamental challenges: the distributional shift between behavior and learned policies in offline RL [24, 25], and the quick adaptation to unseen tasks with minimal data. Existing approaches can be broadly categorized into optimization-based methods like MAML [26] and MACAW [9], and context-based methods such as PEARL [27], VariBAD [28], CORRO [23], CSRO [11], and UNICORN [29]. Recent studies of in-context RL attempt to leverage the in-context learning capability of transformers to improve RL’s generalization via casting RL as an across-episodic sequential prediction problem with few-shot prompting at test time [30, 31], such as AD [32] and DPT [33]. In general, the above approaches rely on high-quality samples or domain knowledge to infer task beliefs during evaluation, limiting their generality with this restricted form of supervision. This inspires us to explore a much broader source of supervision from natural language.
LLMs for RL. Existing studies on leveraging LLMs for RL domains can be categorized into three main threads: LLMs as rewards, LLMs as policies, and language-conditioned policy learning. The first kind uses LLMs or VLMs (vision-language models) to automate the generation and shaping of reward functions [34, 19, 35, 20] or state representations [36], providing denser information to facilitate RL training. The second kind utilizes LLMs or VLMs as the policy backbone [16, 17, 37, 18]. Vision-language-action (VLA) models provide a direct instantiation of using pre-trained VLMs for robotics, fine-tuning visually-conditioned VLMs to generate robot control actions, such as RT-2 [38], OpenVLA [39], and $\pi_{0.5}$ [40]. In contrast, we leverage LLMs to obtain task representations for offline meta-RL and train conventional RL policies based on decision transformer or diffuser architectures.
@@
Language-conditioned policy. We focus on the third kind that learns a language-conditioned policy, extracting the world knowledge encoded in LLMs to help train offline meta-RL agents. Previous studies acquire language embeddings using pre-trained LLMs like BERT and GPT [41], while these embeddings are learned independently from decision tasks and may fail to capture domain-related information. Early methods develop rule-based [42] or task-specific [43] intermediate representations to capture task-related semantics, which can require a laborious design and lack scalability. Under an imitation learning paradigm, BC-Z [21] trains a vision-based robotic manipulation policy conditioned on pre-trained language embeddings, and BAKU [44] produces a simple transformer architecture that fuses multi-modal vision-language information and temporal context for multi-task decision-making. To improve language grounding, [45] translates natural language to task language with a referential game, and [22] uses supervised fine-tuning to enable bidirectional translation between language-described skills and rollout data. Overall, the challenges include the inability to capture environment dynamics, knowledge misalignment due to semantic gaps, and limited model scalability. In the paper, we tackle these challenges to fully harness LLMs for building generalist decision agents.
@@ 3.1 Problem Statement
We consider a language-conditioned offline meta-RL setting. Tasks follow a distribution $M_{k}\!\!=\!\!\langle\mathcal{S},\mathcal{A},\mathcal{T}_{k},\mathcal{R}_{k},\gamma\rangle\!\!\sim\!\!P(M)$, sharing the same state-action spaces while varying in the reward and state transition functions, i.e., environment dynamics. For each training task $k$, our system receives a user-provided natural language supervision $l_{k}$ that describes the task (e.g., open the door), paired with an offline dataset $\mathcal{D}_{k}=\sum_{i}(s_{i}^{k},a_{i}^{k},r_{i}^{k},s_{i}^{{}^{\prime}k})$ collected by arbitrary behavior policies. The agent can only access the offline datasets $\sum_{k}(l_{k},\mathcal{D}_{k})$ to train a generalist policy $\pi(a|s,l)$ to follow language instructions. At test time, natural language is used to reference learned decision perceptions or describe new ones. The agent can perform text-to-decision generation in test environments in a zero-shot manner, given any language instruction $l_{\text{new}}$ at hand. The primary objective is to obtain a highly generalizable policy that achieves strong zero-shot performance on unseen tasks.
@@ 3.2 Dynamics-Aware Decision Embedding
Natural language is a complex and unbounded representation of human instruction, and training policies to harness language supervision is nontrivial. A standard approach is to convert language instructions to simpler latent embeddings using popular pre-trained language models that can draw on a wealth of knowledge learned from copious amounts of text. However, directly applying these models to decision domains could fail to solve user-specific problems, as they are typically learned independently of decision-making tasks. A major challenge is that LLMs trained on text corpora typically lack grounding in the interacting environment and fail to capture its dynamics.
@@
For effective grounding, the agent should comprehend environment dynamics of underlying decision tasks. RL typically learns from active interactions with the outer environment. The world model, i.e., the reward and state transition functions $p(s^{\prime},r|s,a)$, fully characterizes the environment and presents a promising alternative to capturing environment dynamics. Hence, we introduce a generalized world model to encode the multi-task decision data into a dynamics-aware embedding space. We separate the reasoning about the world model into two parts: i) encoding the dynamics-specific information into a latent embedding, and ii) decoding environment dynamics conditioned on that embedding.
@@
First, we use a trajectory encoder $\phi$ to abstract multi-task decision trajectories into a compact embedding $z$ that captures task-specific dynamics. Specifically, we “tokenize" different elements (i.e., state, action, and reward) in the raw sequence $\tau=(s_{0},a_{0},r_{0},...,s_{L},a_{L},r_{L})\sim\mathcal{D}_{k}$, by lifting them to a common representation space using element-specific tokenizers $f_{\phi}^{s},f_{\phi}^{a},f_{\phi}^{r}$ as
$\tau_{e}=(e_{0}^{s},~{}e_{0}^{a},~{}e_{0}^{r},~{}...,~{}e_{L}^{s},~{}e_{L}^{a},~{}e_{L}^{r}),~{}~{}\text{where}~{}e_{t}^{s}\!=\!f_{\phi}^{s}(s_{t}),e_{t}^{a}\!=\!f_{\phi}^{a}(a_{t}),e_{t}^{r}\!=\!f_{\phi}^{r}(r_{t}),~{}\forall t\in[0,L].$ (1)
@@
We use a bi-directional transformer $E_{\phi}$ to extract the dynamics-aware embedding from input tokens as $z_{k}\!=\!E_{\phi}(\tau_{e})$. The bi-directional structure allows for capturing the forward and inverse dynamics.
@@
Second, we introduce a decoder $\varphi$ containing a reward model $\mathcal{R}_{\varphi}$ and state transition model $\mathcal{T}_{\varphi}$. The latent embedding is augmented into the input to predict the instant reward $\hat{r}_{t}\!=\!\mathcal{R}_{\varphi}(s_{t},a_{t};z_{k})$ and next state $\hat{s}_{t+1}\!=\!\mathcal{T}_{\varphi}(s_{t},a_{t};z_{k})$. The encoder-decoder pipeline is jointly trained by minimizing the reward and state transition prediction error conditioned on the decision embedding as
$\mathcal{L(\phi,\varphi)}=\mathbb{E}_{\tau\sim\mathcal{D}_{k}}\left[\mathbb{E}_{z_{k}\sim\phi(\tau)}\left[\mathbb{E}_{t}\left[\left(r_{t}-\mathcal{R}_{\varphi}(s_{t},a_{t};z_{k})\right)^{2}+\left(s_{t+1}-\mathcal{T}_{\varphi}(s_{t},a_{t};z_{k})\right)^{2}\right]\right]\right]\!\!,~{}\forall k.$ (2)
@@
In practice, we randomly sample a trajectory from the dataset as $\tau\!\sim\!\mathcal{D}_{k}$ and use its decision embedding to decode the dynamics of other trajectories in the same dataset as $\tau^{*}\in\mathcal{D}_{k}\backslash\tau$. The reason is that, since the embedding has access to the entire trajectory’s information, using it to decode the same trajectory can lead to deceptive traps during training. After proper training, we freeze the world model for the subsequent learning phases.
@@ 3.3 Contrastive Language-Decision Pre-training
To successfully instruct generalist training with human language, we must bridge the semantic gap between text and the “decision modality". Inspired by CLIP [14], we predict which textual description goes with which training task, bridging their semantic gap via contrastive learning. The core idea is to connect the language supervision with corresponding decision embeddings, distilling the world model structure to the text modality and aligning text embeddings to comprehend environment dynamics.
Formally, given a batch of $N$ (trajectory $\tau$, textual task description $l$) pairs, we predict which of the $N\times N$ possible (trajectory, text) pairings across a batch actually occurred. We use the pre-trained trajectory encoder in Sec. 3.2 to extract the dynamics-aware decision embedding from the trajectory as $z=\phi(\tau)$. The text encoder is a transformer-based model $\psi$ that could be initialized from one of a wide variety of popular language models, such as CLIP [14] or T5 [46]. Then, the text embedding $z_{T}$ is easily derived by tokenizing the textual instruction $l$ and feeding it to the encoder as $z_{T}=\psi(l)$.
@@
$\text{sim}(\tau,l)\!=\!\exp(\alpha)\cdot\frac{\phi(\tau)W_{D}\cdot\psi(l)W_{T}}{||\phi(\tau)W_{D}||\cdot||\psi(l)W_{T}||},~{}~{}\text{sim}(l,\tau)\!=\!\exp(\alpha)\cdot\frac{\psi(l)W_{T}\cdot\phi(\tau)W_{D}}{||\psi(l)W_{T}||\cdot||\phi(\tau)W_{D}||},$ (3)
where $\alpha$ is a learnable temperature controlling the range of similarities, and $W_{D}$ and $W_{T}$ are learnable projections that map feature representations of each modality to a joint multi-modal embedding space. The decision-to-text $p(\tau)$ and text-to-decision $p(l)$ similarity scores in the batch are calculated as
@@
Let $q(\tau)$ and $q(l)$ denote the ground-truth similarity scores, where negative pairs have a probability of 0 and positive pairs have 1. We optimize a symmetric cross-entropy loss over similarity scores as
$\mathcal{L}(\psi)=0.5\cdot\mathbb{E}_{\tau,l}\left[\text{CE}\left(p(\tau),q(\tau)\right)+\text{CE}\left(p(l),q(l)\right)\right],$ (5)
where CE is the cross-entropy between two distributions. To effectively adapt the text encoder to the semantics of decision modality at a lightweight computing cost, we employ low-rank adaptation (LoRA) [47] to fine-tune the text encoder with the above contrastive loss. Experiments in Appendix E present the analysis of its efficiency and superiority compared to full-parameter fine-tuning.